Good day,

For the reaction
3A(g)+2B(g)⇌C(g)
Kc= 72.2 at a temperature of 237 ∘C
Calculate the value of Kp
Express your answer numerically.

Thanks
1
Expert's answer
2020-08-06T03:55:14-0400

Kp = Kc×(R×T)Δn

There is 1 mol of gas on the product side of the chemical equation, and there are 3+2=5 mol of gaseous reactants, so Δn = (1-5) = -4 in this case.

T = 237 + 273.15 = 510.15 K

R = 0.08206 L×atm/mol×K

Kp = Kc×(R×T)Δn = 72.2×(0.08206×510.15)-4 = 72.2×(41.86)-4 = 72.2×3.25×10-7 = 234.65×10-7 = 2.34×10-5

Answer: 2.34×10-5